# Learning from B Cell Evolution: Adaptive Multi-Expert Diffusion for Antibody Design via Online Optimization

### v1

@@ Abstract
Recent advances in diffusion models have shown remarkable potential for antibody design, yet existing approaches apply uniform generation strategies that cannot adapt to each antigen’s unique requirements. Inspired by B cell affinity maturation—where antibodies evolve through multi-objective optimization balancing affinity, stability, and self-avoidance—we propose the first biologically-motivated framework that leverages physics-based domain knowledge within an online meta-learning system. Our method employs multiple specialized experts (van der Waals, molecular recognition, energy balance, and interface geometry) whose parameters evolve during generation based on iterative feedback, mimicking natural antibody refinement cycles. Instead of fixed protocols, this adaptive guidance discovers personalized optimization strategies for each target. Our experiments demonstrate that this approach: Our experiments demonstrate that this approach: (1) discovers optimal SE(3)-equivariant guidance strategies for different antigen classes without pre-training, preserving molecular symmetries throughout optimization; (2) significantly enhances hotspot coverage and interface quality through target-specific adaptation, achieving balanced multi-objective optimization characteristic of therapeutic antibodies; (3) establishes a paradigm for iterative refinement where each antibody-antigen system learns its unique optimization profile through online evaluation; (4) generalizes effectively across diverse design challenges, from small epitopes to large protein interfaces, enabling precision-focused campaigns for individual targets.
@@ Introduction
Computational antibody design remains a fundamental challenge in therapeutic development, requiring simultaneous optimization of hotspot coverage, structural stability, and binding interface quality (Fischman and Ofran 2018; Norman et al. 2020). While recent diffusion-based methods like RFdiffusion (Watson et al. 2023) and RFAntibody (Luo et al. 2022b) show promise for generating novel protein structures, they lack mechanisms to incorporate real-time constraints during generation, often producing designs that fail to meet multiple competing objectives (Eguchi et al. 2022).
Current approaches face three fundamental limitations: (1) they generate structures through unconditional diffusion without target-specific guidance (Watson et al. 2023; Trippe et al. 2023), (2) they cannot balance multiple objectives during generation, requiring inefficient post-hoc filtering (Jin et al. 2022; Shuai, Ruffolo, and Gray 2021), and (3) they either ignore physical constraints or require extensive labeled data to train property predictors (Dauparas et al. 2022; Hsu et al. 2022). These limitations significantly reduce their effectiveness for therapeutic antibody design, where each target presents unique challenges (Akbar et al. 2022; Mason et al. 2021). Prior work confirms that incorporating physical properties improves model performance in biomolecular design (Qiu, Feng, and Poczos 2025; Luo et al. 2022a).
@@
  • Multi-expert guidance system: We develop specialized guidance modules for van der Waals interactions (Jumper et al. 2021), molecular recognition (Gainza et al. 2020), energy balance (Ingraham et al. 2019), and interface geometry (Schneider et al. 2022), each providing targeted gradients during diffusion while maintaining equivariance (Satorras, Hoogeboom, and Welling 2021).
@@
  • Online parameter adaptation: Using Bayesian optimization with Gaussian processes, our framework learns optimal guidance strategies for each antigen-antibody pair through iterative batch evaluation, automatically discovering target-specific temporal profiles without requiring extensive pre-training (Raybould et al. 2019; Shanehsazzadeh et al. 2023a).
Experiments across diverse targets demonstrate substantial improvements: 7% reduction in CDR-H3 RMSD, 9 percentage points increase in hotspot coverage, and consistently enhanced metrics across all evaluation criteria. This balanced optimization addresses the critical ”weakest link” problem in antibody design. By combining physics-based guidance with online learning, our work opens new directions for adaptive approaches in biomolecular design (Ruffolo et al. 2023; Evans et al. 2022).
@@
Recent generative approaches shifted to de novo design. DiffAb (Luo et al. 2022a) introduced diffusion models for joint sequence-structure generation but operates in internal coordinates, limiting inter-chain modeling. RFAntibody (Adolf-Bryfogle, Toth, and Bahl 2024) addressed this using SE(3)-equivariant backbone generation in Cartesian space. However, current methods lack both physicochemical constraint enforcement and target-specific adaptation during generation. This results in high-throughput campaigns producing substantial fractions of structurally unreasonable or experimentally nonviable designs.
@@
Frame Representations and Diffusion Process
@@
The forward diffusion process adds noise over time steps $t\in\{0,1,\ldots,T\}$. Rotations follow the IGSO3 distribution:
[…]
@@
with score function $\nabla_{\mathbf{R}_{t}}\log q(\mathbf{R}_{t}|\mathbf{R}_{0})=\frac{1}{\sigma_{t}^{2}}\text{skew}(\mathbf{R}_{t}^{T}\mathbf{R}_{0})$, where $\text{skew}(\mathbf{M})=\frac{\mathbf{M}-\mathbf{M}^{T}}{2}$.
Translations follow Gaussian diffusion with center-of-mass constraint:
@@
Reverse Diffusion and Network Architecture The reverse process learns to denoise by estimating clean structure $\mathbf{T}_{0}$ from noisy $\mathbf{T}_{t}$:
$p_{\theta}(\mathbf{T}_{t-1}|\mathbf{T}_{t})$ $=q(\mathbf{T}_{t-1}|\mathbf{T}_{t},\hat{\mathbf{T}}_{0}),$ (4)
@@
While RFdiffusion excels at generating diverse protein structures, it faces a critical challenge in antibody design: maintaining balanced performance across multiple essential metrics. In antibody engineering, success requires simultaneous optimization of structural accuracy (CDR-H3 RMSD), prediction confidence (pAE, ipAE), interface quality (shape complementarity, buried surface area), and biophysical properties (VDW interactions, stability).
The fundamental issue is that RFdiffusion’s purely data-driven approach lacks mechanisms to enforce multi-objective balance. This leads to the ”weakest link” problem—a single failing metric renders the entire design experimentally nonviable, regardless of excellence in other areas. In high-throughput campaigns, even if 90% of metrics are excellent, failure in the remaining 10% eliminates the design from experimental consideration.
@@
Energetic Landscape: The binding free energy can be decomposed as:
@@
Shape Complementarity Principle: Following the lock-and-key model extended by induced fit theory, optimal binding requires:
@@
where $S_{\text{buried}}$ represents buried surface area, and $d_{\text{gap}}$ captures interface gaps penalized by decay length $\lambda$.
@@
The temporal activation of experts follows an information-theoretic principle. At timestep $t$, the mutual information between the noisy structure $\mathbf{x}_{t}$ and the target structure $\mathbf{x}_{0}$ is:
@@
This suggests that different structural features become identifiable at different noise levels, justifying our adaptive expert routing based on the signal-to-noise ratio at each timestep.
Convergence Guarantees under Multi-Objective Optimization
Theorem (Pareto Convergence): Under mild assumptions (Lipschitz continuity of expert losses, bounded gradients), our multi-expert guided diffusion converges to a Pareto-optimal solution:
$\begin{split}\mathcal{P}^{*}=\{\mathbf{x}:\nexists\mathbf{x}^{\prime}\text{ s.t. }L_{i}(\mathbf{x}^{\prime})&\leq L_{i}(\mathbf{x})\,\forall i\\ \text{and }L_{j}(\mathbf{x}^{\prime})&<L_{j}(\mathbf{x})\text{ for some }j\}\end{split}$ (9)
This theoretical guarantee ensures that our method finds solutions that cannot be improved in any objective without degrading another, aligning with the multi-criteria nature of therapeutic antibody design.
@@
We present an adaptive physics-guided diffusion framework for antibody design that integrates multi-expert guidance with online learning. Our approach enhances the standard SE(3) equivariant diffusion process by incorporating domain-specific physical constraints through a dynamically weighted ensemble of expert modules. Figure 2 presents our adaptive framework. The system combines physics-based expert guidance (left) with the antibody design pipeline, where guided diffusion generates structures, ProteinMPNN designs sequences, and RoseTTAFold2 validates results. A Bayesian optimization feedback loop (bottom) continuously learns optimal guidance parameters from evaluation metrics.
@@
$\mathbf{X}_{t-1}=\boldsymbol{\mu}_{\theta}(\mathbf{X}_{t},t)+\boldsymbol{\Sigma}_{\theta}^{1/2}(\mathbf{X}_{t},t)\cdot\mathbf{z}-\lambda(t)\cdot\mathbf{g}(\mathbf{X}_{t},t)$ (10)
where $\boldsymbol{\mu}_{\theta}$ and $\boldsymbol{\Sigma}_{\theta}$ are the predicted mean and covariance from the denoising network, $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, and $\mathbf{g}(\mathbf{X}_{t},t)$ represents the combined guidance gradient:
$\mathbf{g}(\mathbf{X}_{t},t)=\sum_{i=1}^{4}w_{i}(t)\cdot\nabla\mathcal{L}_{i}(\mathbf{X}_{t})$ (11)
@@
where $\lambda_{\text{base},i}$ is the base strength for expert $i$, and $f_{\text{temporal}}(t,\alpha,\beta)$ is the temporal modulation function based on the Beta distribution:
@@
where $t_{\text{norm}}=(t-1)/(T-1)\in[0,1]$ is the normalized timestep, $B(\cdot;\alpha,\beta)$ denotes the Beta probability density function, and $t_{\text{mode}}=(\alpha-1)/(\alpha+\beta-2)$ for $\alpha,\beta>1$ represents the mode of the distribution. The peak strength factor $\lambda_{\text{peak}}$ scales the maximum activation, while the shape parameters $\alpha$ and $\beta$ control the temporal profile, enabling diverse activation patterns from early-peaking to late-peaking distributions.
Figure 2: Adaptive multi-expert diffusion framework with online learning for antibody design.
@@
This expert ensures proper CDR coverage of epitope hotspots. For each hotspot $h\in\mathcal{H}$, we encourage proximity to CDR residues:
@@
where $n_{c}$ denotes the number of inter-molecular contacts within threshold $d_{c}$, $[\tau^{-},\tau^{+}]$ defines the target range, and $\phi^{-}(\cdot)$, $\phi^{+}(\cdot)$ are penalty functions for sparse and dense packing, respectively. The gradient modulates CDR positioning to optimize interface contacts.
@@
Our routing mechanism dynamically activates experts based on real-time structural analysis (Figure 3), ensuring that computational resources focus on the most critical problems while maintaining SE(3)-equivariance. The system computes problem severity scores:
@@
where $f_{i}$ evaluates current structural metrics to determine the severity of problems relevant to expert $i$.
@@
Figure 3: Dynamic expert activation across diffusion timesteps. Colored regions indicate when each expert is active based on detected structural problems, demonstrating the adaptive nature of our routing mechanism.
@@
$\nabla\mathcal{L}_{i}(\mathbf{R}\mathbf{X}+\mathbf{t})=\mathbf{R}\nabla\mathcal{L}_{i}(\mathbf{X})$ (24)
for any rotation $\mathbf{R}\in\text{SO}(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$. This is achieved by basing all computations on pairwise distances, relative orientations, and other invariant descriptors rather than absolute coordinates. The weighted combination of equivariant gradients maintains equivariance: $\mathbf{g}(\mathbf{X}_{t},t)=\sum_{i}w_{i}(t)\cdot\nabla\mathcal{L}_{i}(\mathbf{X}_{t})$ remains equivariant since weights $w_{i}(t)$ are scalar functions of invariant metrics.
@@
We employ Bayesian optimization with Gaussian processes (GP) to adaptively learn the optimal Beta distribution parameters for each antibody-antigen system. Our approach models a two-dimensional parameter space $\boldsymbol{\theta}=(\alpha,\beta)\in\Theta$, where $\alpha$ and $\beta$ are the shape parameters controlling the temporal modulation profile.
The optimization process maintains a probabilistic surrogate model of the objective function using a Gaussian process:
@@
where the metrics include CDR-H3 backbone RMSD, predicted aligned error (pAE), and interface pAE (ipAE), each normalized by appropriate constants $\nu_{m}$.
@@
To handle stochastic evaluations, we aggregate results from similar parameters to reduce noise. The framework periodically re-evaluates promising configurations to obtain more reliable estimates. This adaptive approach automatically discovers case-specific Beta distribution profiles—from early-peaking ($\alpha<\beta$) for global structure guidance to late-peaking ($\alpha>\beta$) for atomic refinement—without manual parameter tuning, typically converging within a reasonable number of evaluations.
@@
We evaluated our adaptive physics-guided framework on six targets from the IgDesign benchmark (Shanehsazzadeh et al. 2023b): IL17A (PDB: 6PPG), ACVR2B (PDB: 5NGV), FXI (PDB: 6HHC), TSLP (PDB: 5J13), IL36R (PDB: 6U6U), and TNFRSF9 (PDB: 6A3W). These targets span 17-40 kDa, cover inflammation/thrombosis/cancer applications, include 1,243 SPR-validated designs, and exhibit diverse binding modes. We annotated 5 hotspot residues per target and generated 1,000 designs each.
@@
Most recent methods are unavailable due to lack of publicly accessible code. We selected: DiffAb (Luo et al. 2022c): torsion-space CDR generation; RFAntibody (Bennett et al. 2024): the current state-of-the-art method without physics guidance, widely recognized for its exceptional performance across diverse antibody design tasks. All used identical inputs: target+hotspots, Trastuzumab framework.
@@
All experiments used 4 NVIDIA RTX A6000 GPUs with the following settings:
@@
Successful designs must satisfy: CDR-H3 RMSD $<$ 3.0 Å, mean pAE $<$ 10.0, and mean interaction pAE (ipAE) $<$ 10.0. This multi-metric evaluation ensures structural accuracy and binding confidence.
@@
Figure 5: Adaptive guidance improves antibody design quality across multiple metrics. Box plots compare baseline RFAntibody (gray) with our guided approach (pink) on six targets (5NGV, 6U6U, 6HHC, 6A3W, 6PPG, 5J13) for (a) CDR-H3 RMSD, (b) interface pAE, (c) mean pAE, and (d) VDW clashing.
@@
Method Success(%) H3-RMSD(Å) pAE iPAE
DiffAb 11.13$\pm$16.95 9.12 $\pm$ 10.08 9.38$\pm$1.16 10.09 $\pm$2.21
RFAntibody 36.67$\pm$14.01 2.08$\pm$1.81 7.13$\pm$0.46 8.49$\pm$1.71
Ours 41.07$\pm$10.51 1.94$\pm$1.36 6.85$\pm$0.39 8.12$\pm$1.31
Method SC BSA BB VDW(kcal/mol) pLDDT
DiffAb 0.552$\pm$0.142 5277$\pm$2341 -15.21$\pm$9.67 0.886$\pm$0.01
RFAntibody 0.549$\pm$0.026 3583$\pm$888 -1.18$\pm$1.27 0.896$\pm$0.011
Ours 0.576$\pm$0.024 3870$\pm$780 -1.33$\pm$0.95 0.893$\pm$0.008
Method Hotspot(%) CDR Inter.(%) Sequence Diversity Electrostatic Energy
DiffAb 16.20$\pm$5.54 0.00$\pm$0.00 0.464$\pm$0.080 -6.26$\pm$11.14
RFAntibody 48.5$\pm$18.1 54.7$\pm$13.55 0.534$\pm$0.004 -3.96$\pm$7.07
Ours 57.5$\pm$12.6 61.3$\pm$11.2 0.532$\pm$0.004 -5.64$\pm$6.10
Ablation Studies
@@
Configuration Success(%) H3-RMSD(Å) pAE iPAE
RFAntibody (Baseline) 36.7$\pm$14.0 2.11$\pm$1.81 7.13$\pm$0.46 8.49$\pm$1.71
+ Multi-Expert Guidance 38.9$\pm$12.5 2.01$\pm$1.58 6.95$\pm$0.42 8.28$\pm$1.65
+ Online Learning (Full) 41.1$\pm$10.5 1.94$\pm$1.36 6.85$\pm$0.39 8.12$\pm$1.31
Configuration SC BSA BB VDW(kcal/mol) pLDDT
RFAntibody (Baseline) 0.549$\pm$0.026 3583$\pm$888 -1.18$\pm$1.27 0.896$\pm$0.011
+ Multi-Expert Guidance 0.562$\pm$0.026 3712$\pm$885 -1.25$\pm$1.15 0.885$\pm$0.014
+ Online Learning (Full) 0.576$\pm$0.024 3870$\pm$780 -1.33$\pm$0.95 0.893$\pm$0.008
Configuration Hotspot(%) CDR Inter.(%) Sequence Diversity Electrostatic Energy
RFAntibody (Baseline) 48.5$\pm$18.1 54.71$\pm$13.55 0.534$\pm$0.004 -3.96$\pm$7.07
+ Multi-Expert Guidance 55.6$\pm$15.2 60.8$\pm$13.5 0.532$\pm$0.005 -5.31$\pm$8.1
+ Online Learning (Full) 57.5$\pm$12.6 61.3$\pm$11.2 0.532$\pm$0.004 -5.64$\pm$6.1
@@
Target Type Optimized Beta Parameters Key Expert Emphasis
$\alpha$ $\beta$
Small epitopes (IL17A, IL36R) 3.2$\pm$0.4 1.8$\pm$0.3 VDW + Recognition
Large interfaces (ACVR2B, TNFRSF9) 2.1$\pm$0.3 3.5$\pm$0.5 Interface + Energy
Buried epitopes (FXI) 2.8$\pm$0.4 2.2$\pm$0.3 Recognition + Energy
Flexible targets (TSLP) 2.0$\pm$0.5 2.0$\pm$0.5 All experts balanced

### v2

@@ Abstract
Recent advances in diffusion models have shown remarkable potential for antibody design, yet existing approaches apply uniform generation strategies that cannot adapt to each antigen’s unique requirements. Inspired by B cell affinity maturation—where antibodies evolve through multi-objective optimization balancing affinity, stability, and self-avoidance—we propose the first biologically-motivated framework that leverages physics-based domain knowledge within an online meta-learning system. Our method employs multiple specialized experts (van der Waals, molecular recognition, energy balance, and interface geometry) whose parameters evolve during generation based on iterative feedback, mimicking natural antibody refinement cycles. Instead of fixed protocols, this adaptive guidance discovers personalized optimization strategies for each target. Our experiments demonstrate that this approach: (1) discovers optimal SE(3)-equivariant guidance strategies for different antigen classes without pre-training, preserving molecular symmetries throughout optimization; (2) significantly enhances hotspot coverage and interface quality through target-specific adaptation, achieving balanced multi-objective optimization characteristic of therapeutic antibodies; (3) establishes a paradigm for iterative refinement where each antibody-antigen system learns its unique optimization profile through online evaluation; (4) generalizes effectively across diverse design challenges, from small epitopes to large protein interfaces, enabling precision-focused campaigns for individual targets.
@@ Introduction
Computational antibody design remains a fundamental challenge in therapeutic development, requiring simultaneous optimization of hotspot coverage, structural stability, and binding interface quality (Fischman and Ofran 2018; Norman et al. 2020). While recent diffusion-based methods like RFdiffusion (Watson et al. 2023) and RFAntibody (Luo et al. 2022) show promise for generating novel protein structures, they lack mechanisms to incorporate real-time constraints during generation, often producing designs that fail to meet multiple competing objectives (Eguchi et al. 2022).
Current approaches face three fundamental limitations: (1) they generate structures through diffusion without target-specific guidance (Watson et al. 2023; Trippe et al. 2023), (2) they cannot balance multiple objectives during generation, requiring inefficient post-hoc filtering (Jin et al. 2022; Shuai, Ruffolo, and Gray 2021), and (3) they either ignore physical constraints or require extensive labeled data to train property predictors (Dauparas et al. 2022; Hsu et al. 2022). These limitations significantly reduce their effectiveness for therapeutic antibody design, where each target presents unique challenges (Akbar et al. 2022; Mason et al. 2021).
@@
  • Multi-expert guidance system: We develop specialized guidance modules for van der Waals interactions (Jumper et al. 2021), molecular recognition (Gainza et al. 2020), energy balance (Ingraham et al. 2019), and interface geometry (Schneider et al. 2022), each providing targeted gradients during diffusion while maintaining equivariance (Satorras, Hoogeboom, and Welling 2021a).
@@
  • Online parameter adaptation: Using Bayesian optimization with Gaussian processes, our framework learns optimal guidance strategies for each antigen-antibody pair through iterative batch evaluation, automatically discovering target-specific temporal profiles without requiring pre-training (Raybould et al. 2019; Shanehsazzadeh et al. 2023).
Experiments across diverse targets demonstrate substantial improvements: 7% reduction in CDR-H3 RMSD, 9% increase in hotspot coverage, 12% better interface pAE, and 5% higher shape complementarity, with enhanced metrics across all evaluation criteria. This balanced optimization addresses the critical “weakest link” problem (Raybould et al. 2019) in antibody design. By combining physics-based guidance with online learning, our work opens new directions for adaptive approaches in biomolecular design.
@@
Recent generative approaches shifted to de novo design. DiffAb (Luo et al. 2022) introduced diffusion models for joint sequence-structure generation but operates in internal coordinates, limiting inter-chain modeling. RFAntibody (Adolf-Bryfogle, Toth, and Bahl 2024) addressed this using SE(3)-equivariant backbone generation in Cartesian space. However, current methods lack both physicochemical constraint enforcement and target-specific adaptation during generation. This results in high-throughput campaigns producing substantial fractions of structurally unreasonable or experimentally nonviable designs (Shen et al. 2024). Additional related work is discussed in Appendix A due to space limitations.
@@
Frame Representations and Forward Diffusion Process
@@
The forward diffusion process adds noise over time steps $t\in\{0,1,\ldots,T\}$. Rotations follow the IGSO3 distribution, $q(\mathbf{R}_{t}|\mathbf{R}_{0})=\text{IGSO3}(\mathbf{R}_{t};\mathbf{R}_{0},\sigma_{t})$ with density:
@@
where $\sigma_{t}$ is the noise level at timestep $t$, controlling the variance of the distribution(Yim et al. 2023). The derivation of the score function and its properties on the SO(3) manifold are detailed in Appendix B. Translations follow Gaussian diffusion with center-of-mass constraint:
@@
Reverse Diffusion and Network Architecture The reverse process learns to denoise the forward diffusion process by estimating clean structure $\mathbf{T}_{0}$ from noisy $\mathbf{T}_{t}$ (Ho, Jain, and Abbeel 2020):
$p_{\theta}(\mathbf{T}_{t-1}|\mathbf{T}_{t})$ $=p(\mathbf{T}_{t-1}|\mathbf{T}_{t},\hat{\mathbf{T}}_{0}),$ (3)
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ is the cumulative product of noise schedule parameters, and $\epsilon_{\theta}(\mathbf{T}_{t},t)$ is the neural network with parameters $\theta$ that predicts the noise added at timestep $t$.
@@
While state-of-the-art generative models excels at generating diverse protein structures, it faces a critical challenge in antibody design: maintaining balanced performance across multiple essential metrics. In antibody engineering, success requires simultaneous optimization of structural accuracy (CDR-H3 RMSD), prediction confidence (pAE, ipAE), interface quality (shape complementarity, buried surface area), and biophysical properties (VDW interactions, stability) (Raybould et al. 2019; Norman et al. 2020; Choi et al. 2018).
The fundamental issue is that RFdiffusion’s purely data-driven approach lacks mechanisms to enforce multi-objective balance. This leads to the ”weakest link” problem—a single failing metric renders the entire design experimentally nonviable (Raybould et al. 2019; Choi et al. 2018), regardless of excellence in other areas. In high-throughput campaigns, even if 90% of metrics are excellent, failure in the remaining 10% eliminates the design from experimental consideration.
@@
Energetic Landscape: The binding free energy ($\Delta G_{\text{bind}}$) can be decomposed as  (Gilson and Zhou 2007):
@@
Shape Complementarity Principle: Following the lock-and-key model extended by induced fit theory (Fischer 1894; Koshland 1958), optimal binding requires:
@@
where Ab and Ag denote antibody and antigen respectively, $S_{\text{buried}}$ represents buried surface area upon complex formation, $S_{\text{total}}$ is the total surface area of each protein, and $d_{\text{gap}}$ captures interface gaps penalized by decay length $\lambda$.
@@
Our adaptive expert routing is motivated by information theory. At timestep $t$, the mutual information between the noisy structure $\mathbf{x}_{t}$ and the target structure $\mathbf{x}_{0}$ is (Austin et al. 2021):
@@
where $\alpha_{t}$ is the signal retention coefficient at timestep $t$. The derivation of this relationship is provided in Appendix C.
This suggests that different structural features become identifiable at different noise levels, justifying our adaptive expert routing based on the signal-to-noise ratio at each timestep. The connection between SNR and feature visibility is detailed in Appendix D.
@@
We present an adaptive physics-guided diffusion framework for antibody design that integrates multi-expert guidance with online learning. Our approach enhances the standard SE(3) equivariant diffusion process (Hoogeboom et al. 2022; Watson et al. 2023) by incorporating domain-specific physical constraints through a dynamically weighted ensemble of expert modules. Figure 2 presents our adaptive framework. The system combines physics-based expert guidance (left) with the antibody design pipeline, where guided diffusion generates structures, ProteinMPNN designs sequences, and RoseTTAFold2 validates results. A Bayesian optimization feedback loop (bottom) continuously learns optimal guidance parameters from evaluation metrics.
@@
$\mathbf{T}_{t-1}=\boldsymbol{\mu}_{\theta}(\mathbf{T}_{t},t)+\boldsymbol{\Sigma}_{\theta}^{1/2}(\mathbf{T}_{t},t)\cdot\mathbf{z}-\lambda(t)\cdot\mathbf{g}(\mathbf{T}_{t},t)$ (8)
where $\boldsymbol{\mu}_{\theta}$ and $\boldsymbol{\Sigma}_{\theta}$ are the predicted mean and covariance from the denoising network, $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, and $\mathbf{g}(\mathbf{T}_{t},t)$ represents the combined guidance gradient:
$\mathbf{g}(\mathbf{T}_{t},t)=\sum_{i=1}^{4}w_{i}(t)\cdot\lambda_{i}(t)\cdot\nabla\mathcal{L}_{i}(\mathbf{T}_{t})$ (9)
where $w_{i}(t)$ are adaptive weights for each expert and $\nabla\mathcal{L}_{i}$ are gradients of expert-specific physics-based loss functions.
@@
where $\lambda_{\text{base},i}$ is the base strength for expert $i$, and $f_{\text{temporal}}(t,\alpha,\beta)$ controls when each expert is most active during denoising. The shape parameters $\alpha$ and $\beta$ are learned via Bayesian optimization to discover optimal temporal profiles—early-peaking ($\alpha<\beta$) for global structure experts, late-peaking ($\alpha>\beta$) for atomic refinement:
@@
where $t_{\text{norm}}=(t-1)/(T-1)$ normalizes the timestep to $[0,1]$, and $B(\cdot;\alpha,\beta)$ is the Beta density function. The shape parameters $\alpha$ and $\beta$ are learned via Bayesian optimization to discover optimal activation timing for each target. Detailed profiles are shown in Appendix E.
Figure 2: Adaptive multi-expert diffusion framework with online learning for antibody design. Due to space constraints, detailed descriptions of each module and their input/output dimensions are provided in Appendix F.
@@
This expert ensures proper CDR coverage of epitope hotspots. For each hotspot $h\in\mathcal{H}$, we encourage proximity to CDR residues(see Appendix G for hotspot definition and selection methodology):
@@
where $n_{c}$ denotes the number of inter-molecular contacts within threshold $d_{c}$, $\tau^{-},\tau^{+}]$ defines the target range, and $\phi^{-}(\cdot)$, $\phi^{+}(\cdot)$ are penalty functions for sparse and dense packing, respectively. The gradient modulates CDR positioning to optimize interface contacts.
@@
Our routing mechanism dynamically activates experts based on real-time structural analysis, ensuring that computational resources focus on the most critical problems while maintaining SE(3)-equivariance. The system computes problem severity scores:
@@
where $f_{i}$ evaluates current structural metrics to determine the severity of problems relevant to expert $i$ (detailed in the Appendix H).
@@
$\nabla\mathcal{L}_{i}(\mathbf{R}\mathbf{T}+\mathbf{t})=\mathbf{R}\nabla\mathcal{L}_{i}(\mathbf{T})$ (22)
for any rotation $\mathbf{R}\in\text{SO}(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$. This is achieved by basing all computations on pairwise distances, relative orientations, and other invariant descriptors rather than absolute coordinates. The weighted combination of equivariant gradients maintains equivariance: $\mathbf{g}(\mathbf{T}_{t},t)=\sum_{i}w_{i}(t)\cdot\nabla\mathcal{L}_{i}(\mathbf{T}_{t})$ remains equivariant since weights $w_{i}(t)$ are scalar functions of invariant metrics.
@@
We employ Bayesian optimization with Gaussian processes (GP) to adaptively learn the optimal Beta distribution parameters for each antibody-antigen system post-generation. Our approach models a two-dimensional parameter space $\boldsymbol{\theta}=(\alpha,\beta)\in\Theta$, where $\alpha$ and $\beta$ are the shape parameters controlling the temporal modulation profile. The optimization process maintains a probabilistic surrogate model of the objective function using a Gaussian process:
@@
where the metrics include CDR-H3 backbone RMSD, predicted aligned error (pAE), and interaction pAE (ipAE), each normalized by appropriate constants $\nu_{m}$.
@@
To handle stochastic evaluations, we aggregate results from similar parameters to reduce noise and periodically re-evaluate promising configurations. This approach automatically discovers optimal Beta parameters for each target without manual tuning. Due to space constraints, convergence analysis and computational requirements are detailed in Appendix I.
@@
We evaluated our framework on six antibody-antigen pairs: IL17A (PDB: 6PPG), ACVR2B (PDB: 5NGV), FXI (PDB: 6HHC), TSLP (PDB: 5J13), IL36R (PDB: 6U6U), and TNFRSF9 (PDB: 6A3W). Prior work including RFAntibody (Bennett et al. 2024) typically evaluated four targets; we expanded this to six therapeutically relevant targets spanning inflammation, thrombosis, and cancer applications (17-40 kDa). This extended benchmark from IgDesign (Shanehsazzadeh et al. 2023) includes 1,243 SPR-validated designs across diverse binding modes. We annotated 5 hotspot residues per target and generated 1,000 designs each, following the field’s established protocol for rigorous benchmarking of antibody design methods.
@@
Most recent methods are unavailable due to lack of publicly accessible code. We selected: DiffAb (Luo et al. 2022): torsion-space CDR generation; RFAntibody (Bennett et al. 2024): the current state-of-the-art method without physics guidance, widely recognized for its exceptional performance across diverse antibody design tasks. For fair comparison, all methods used identical inputs (target structure, hotspot annotations, and Trastuzumab framework).
@@
All experiments used 4 NVIDIA RTX A6000 GPUs with the following settings in Table 1.
@@
Successful designs must satisfy: CDR-H3 RMSD << 3.0 Å, mean pAE << 10.0, and mean interaction pAE (ipAE) << 10.0. This multi-metric evaluation ensures structural accuracy and binding confidence.
@@
Figure 4: Adaptive guidance improves antibody design quality across multiple metrics. Box plots compare baseline RFAntibody (gray) with our guided approach (pink) on six targets for (a) CDR-H3 RMSD, (b) interaction pAE, (c) mean pAE, and (d) NMA stability.
Method Success(%) H3-RMSD(Å) pAE iPAE
[…]
Method SC BSA VDW(kcal/mol) pLDDT
[…]
Method Hotspot(%) CDR Int.(%) Seq. Div. Elec. Energy
[…]
@@
Config Success H3-RMSD pAE iPAE
[…]
Config SC BSA VDW pLDDT
[…]
Config Hotspot CDR Seq.D Elec.E
[…]
@@
Target Type Beta Params Key Expert
$\alpha$ $\beta$ Emphasis
[…]
(IL17A, IL36R)
[…]
(ACVR2B, TNFRSF9) Energy
[…]
(FXI) Energy
[…]
(TSLP) balanced